Encyclopedia Astronautica
2008.04.26 - ISS On-Orbit Status 04/26/08


Saturday - half-day off for CDR Volkov, FE-1 Kononenko and FE-2 Reisman.

The crew performed the regular weekly three-hour task of thorough station cleaning. ("Uborka", usually done on Saturdays, includes removal of food waste products, cleaning of compartments with vacuum cleaner, damp cleaning of the Service Module (SM) dining table, other frequently touched surfaces and surfaces where trash is collected, as well as the FE's sleep station with a standard cleaning solution; also, fan screens and grilles are cleaned to avoid temperature rises. Special cleaning is also done every 90 days on the HEPA (high-efficiency particulate air) bacteria filters in the Lab.)

As part of the house cleaning, CDR Volkov conducted preventive maintenance cleaning on the VPkhO & VPrK fan grilles in the SM, V3 screen in the DC1 Docking Module and TsV2 grille in the FGB (Funktsionalnyi-Grusovoi Blok).

The CDR later completed the routine servicing of the SOZh system (Environment Control & Life Support System, ECLSS) in the SM including ASU toilet facilities systems/replaceables and today also the periodic cleaning of the pre-filter of the POTOK air filtration system.

At ~9:15am EDT, the crewmembers held their regular WPC (Weekly Planning Conference) with the ground, discussing next week's "Look-Ahead Plan" (prepared jointly by MCC-H and TsUP/Moscow timeline planners) via S-band/audio, reviewing the monthly calendar, upcoming activities, and any concerns about future on-orbit events.

After Houston flight controllers deactivated the CDRA (Carbon Dioxide Removal Assembly) yesterday and cooling is no longer required, Flight Engineer Reisman today disconnected the ITCS LTL QD (Internal Thermal Control System/Low Temperature Loop/Quick Disconnect) jumper to the CDRA rack (at loc. LAB1D6).

For his VolSci (Voluntary Science) program today, Garrett reviewed OBT (Onboard Training) material for the SLAMMD (Space Linear Acceleration Mass Measurement Device) equipment on the HRF1 (Human Research Facility 1) rack, set up the equipment and performed a configuration check (control runs with no assistance) of the hardware to verify calibration & communication for upcoming experiment ops. (As 'Operator', Garrett will be the safety 'spotter' for SLAMMD. The experiment provides an accurate means of determining the on-orbit mass of humans spanning the range from the 5th percentile Japanese female and the 95th percentile American male. The procedure, in accordance with Newton's 2nd Law of Motion, finds the mass by dividing force, generated by two springs inside the SLAMMD drawer, by acceleration measured with a precise optical instrument that detects the position versus time trajectory of the SLAMMD guide arm and a micro controller which collects the raw data and provides the precise timing. The final computation is done via portable laptop computer with SLAMMD unique software. To calculate their mass, crewmembers wrap their legs around a leg support assembly, align the stomach against a belly pad and either rest the head or chin on a head rest. For calibration, an 18-lbs. mass is used at different lengths from the pivot point, to simulate different mass values. Allowable crew mass range is from 90 to 240 lbs. During the SLAMMD control run, cables or other crew members must be out of the way so the hardware moves freely.)

The FE-2 finished the first session for Inc-17 (and the 28th aboard ISS) of the experiment EarthKAM (Earth Knowledge Acquired by Middle School Students) from the University of California in San Diego, by disassembling and stowing the hardware. (EarthKAM was activated on 4/19, for the first time in Node-2, at the nadir hatch window, since the Lab science window still needs to be shuttered to protect it against outgassing from the SPDM (Special Purpose Dexterous Manipulator) Dextre. EarthKAM is an education program that enables thousands of students to photograph and examine Earth from the unique perspective of space, integrating the excitement of ISS with middle-school education. The student requests are uplinked in a camera control file to an A31p SSC (Station Support Computer) laptop which then activates the camera, a DCS 760 electronic still camera with 50mm (f/1.4) lens, at specified times and receives the digital images from the camera's storage card on its hard drive, for subsequent downlink via OPS LAN. EarthKAM captured 1033 images for the data-gathering session from 4/22 to 4/26. During this go-around, 60 schools participated (approximately 2800 students) from Argentina, Canada, Columbia, Germany, Japan, Mexico, Spain, South Korea and the United States. 681 images were taken with the 50mm lens and 352 were captured with the 180mm lens after Garrett's lens change on 4/24.)

Afterwards, Reisman returned the two stowage bags #1007 & 1009 from their temporary location at NOD2P2 (to make room for EarthKAM) to the NOD2D2 nadir hatch stowage location.

Volkov & Kononenko had their weekly PFCs (Private Family Conferences) via S-band/audio and Ku-band/MS-NetMeeting application (which displays the uplinked ground video on the SSC-9 laptop), Sergey at 7:10am, Oleg at 8:40am.

The crew completed their regular 2.5-hr. physical workout program (about half of which is used for setup & post-exercise personal hygiene) on the CEVIS cycle ergometer (FE-2), TVIS treadmill (CDR, FE-1), RED resistive exercise device (FE-2) and VELO bike with bungee cord load trainer (CDR, FE-1).

Afterwards, Garrett downloaded the crew's exercise data file to the MEC (Medical Equipment Computer) for downlink, as well as the daily wristband HRM (Heart Rate Monitor) data of the workouts on RED, followed by their erasure on the HRM storage medium (done six times a week).

Working from the Russian voluntary 'time permitting' task list, Kononenko performed a session of the Russian "Uragan" (hurricane) earth-imaging program, using the Nikon D2X digital camera with 300-800 mm telephoto lens. (Targets uplinked for today were Ugra National Park, the Don river valley, general views of the Southern Urals & the Volga-Akhtubinsk river flood plain from Volgograd to the Caspian Sea, water flooding followed hydropower plant gate opening, oil spills on the Caspian Sea, overall views of the Aral Sea with dust storms (if visible), the tips of Pamir glacier tongues (in nadir), mountain peaks of Tibet with glaciers, major peaks of the Himalayas, the Dniestr River valley, the Kerch Strait, Mzyt river valley, the site of future Olympic games from Adler airport to the upper reach of the river, glaciers on the north slope of the Greater Caucasus Range, the Kolka glacier, and general views of the Caspian Sea).

Also from the discretionary task list, Oleg conducted a KPT-3 session to make observations and take aerial KPT-3 photography of water contamination in the Kerch Strait for Russia's Environmental Safety Agency (ECON) using the Nikon D2X digital camera with SIGMA 300-800mm telephoto lens.

Weekly Science Update (Expedition Seventeen -- Week 1)

ALTCRISS (Alteino Long Term monitoring of Cosmic Rays on the ISS): Measurements continue in FGB module. Inc16 science data returned on 15S.

ANITA: Completed.

BCAT-3 (Binary Colloidal Alloy Test 3): Reserve.

CARDIOCOG-2: Completed.

CCISS (Cardiovascular & Cerebrovascular Control on Return from ISS): Reserve.

CFE (Capillary Flow Experiment): Reserve.

CW/CR (Cell Wall/Resist Wall) in EMCS (European Modular Cultivation System): The Cell Wall/Resist Wall Experiment Containers are located in EMCS. Water hydration troubleshooting is ongoing. B3 and B4 hydration succeeded.

CSI-2/CGBA (CGBA Science Insert #2/Commercial Generic Bioprocessing Apparatus): In progress.

CGBA-2 (Commercial Generic Bioprocessing Apparatus 2): Complete.

CSLM-2 (Coarsening in Solid-Liquid Mixtures 2): In progress.

EarthKAM (Earth Knowledge Acquired by Middle School Students): EarthKAM captured 1033 images for the data-gathering session from 4/22 to 4/26. Sixty schools participated (approximately 2800 students) from schools in Argentina, Canada, Columbia, Germany, Japan, Mexico, Spain, South Korea and the United States. 681 images were taken with the 50mm lens and 352 were captured with the 180mm lens. The camera was setup in the Node-2 nadir hatch window.

ELITE-S2 (Elaboratore Immagini Televisive - Space 2): Planned.

EPO (Educational Payload Operations): Reserve.

ETD (Eye Tracking Device): Inc16 science data were returned on 15S.

EuTEF (European Technology Exposure Facility): On 4/20, EuTEF platform encountered a MIL Bus error. Ground operators performed a graceful shutdown, power cycled and fully recovered the platform and its instruments. - DEBIE-2: Link error still under investigation; - DOSTEL: On-going science acquisition; - EuTEMP: Currently inactive as planned; - EVC: Further troubleshooting is on-going; - EXPOSE: On-going science acquisition; - FIPEX: Long-duration measurements with sensors RAM1, RAM4 and ZENITH8 have started on 4/23; - MEDET: On-going science acquisition, running nominal, except sample wheel of the Spectrometer is blocked on "park" position; - PLEGPAY: Langmuir Probe run was successfully performed during the 15S undocking on 4/19 via direct commanding; - TRIBOLAB: On 4/15, the Pin On Disk (POD1) run was suddenly stopped after we encountered much higher than anticipated friction coefficients measurements. After analysis, it was decided to resume this POD1 run on 4/21. But TRIBOLAB went again immediately to a stop mode again, meaning that POD1 run is considered completed.

FSL (Fluid Science Laboratory): The mechanical troubleshooting of the FSL Drawers was successful on 4/23, as FSL activation from ground on 4/24 confirmed that all the FSL laser switches are properly engaged and fully functional.

GEOFLOW: Start of GEOFLOW is pending further FSL commissioning activities (check of FSL optical modes) currently planned for next week.

IMMUNO (Neuroendocrine & Immune Responses in Humans During & After Long Term Stay at ISS): Inc16 Blood and urine samples have been returned with 15S and handed over to science team.

InSPACE-2 (Investigating the Structure of Paramagnetic Aggregates from Colloidal Emulsions 2): In progress.

Integrated Immune: In progress.

KUBIK-FM1/ KUBIK-FM2 Centrifuge/Incubators: Completed.

LOCAD-PTS (Lab-on-a-Chip Application Development-Portable Test System): Complete.

MISSE (Materials ISS Experiment): Ongoing.

MTR-2 (Russian radiation measurements): Passive dosimeters measurements in DC1 'Pirs'.

MULTIGEN-1: Completed.

MSG-SAME (Microgravity Science Glovebox): Complete.

NOA-2 (Nitric Oxide Analyzer): Planned.

NUTRITION/REPOSITORY: In progress.

PMDIS (Perceptual Motor Deficits in Space): Complete.

SAMS/MAMS (Space & Microgravity Acceleration Measurement Systems): Ongoing.

SAMPLE: Inc16 samples have been downloaded with 15S and handed over to science team.

SLEEP (Sleep-Wake Actigraphy & Light Exposure during Spaceflight): In progress.

SOLAR (Solar Monitoring Observatory): The Sun observation is now closed (Beta angles higher than 24 deg), and will resume on 4/30. On 4/18, the instruments were put in safe mode for the 15S undocking. - SOVIM: Waiting for the Sun; - SOLSPEC: Waiting for the Sun; - SOLACES: Waiting for the Sun.

SPHERES (Synchronized Position Hold, Engage, Reorient, Experimental Satellite): In progress.

Swab (Characterization of Microorganisms & Allergens in Spacecraft): Complete.

TRAC (Test of Reaction & Adaptation Capabilities): Planned.

ULTRASOUND: Planned.

WAICO #1 (Waving and Coiling of Arabidopsis Roots at Different g-levels): The 4 ECs (Experiment Containers) of Rotor B are yet to be retrieved by the crew from the blocked Rotor B.

CEO (Crew Earth Observations): 'Greetings to the Increment 17 crew from the staff of Crew Earth Observations! We look forward continuing to work with you in support of our payload. We have already started to receive your imagery; through 4/20 we have received 73 frames including an excellent, unsolicited sequence for one of our science targets: Santorini Volcanic Complex, Greece. We will evaluate this to decide whether any additional imagery of this target is required. By way of continuity we have received a total 18,678 frames through 4/18 for Increment 16. This included excellent 400mm views of the Slate Island Impact Crater site in Lake Superior. Only 800mm views are needed to complete this site. Nice going! Also your striking view of the Cerro Culiacan volcano in west-central Mexico will be published on NASA/GSFC's Earth Observatory website this weekend. It is a fine example of a radial drainage pattern and the integration of agriculture in rich volcanic soils. Good shot!'

CEO photo targets uplinked for today were Chongqing, China (nadir pass over this famous city situated on the Yellow River in the Red Basin. The population of Chongqing proper is 4 million, but 31 million cluster in the county of Chongqing. Images looking obliquely at the mountain walls around the basin in which Chongqing lies can also reveal the density and altitude of the persistent smog blanket), Dhaka, Bangladesh (the metro area population is almost 12 million. The crew was to shoot margins of the city. Looking left, two famous rivers were the visual cue: Dhaka lies on the far side of the Ganges River, between the very wide Brahmaputra River and the much smaller Meghna River), and Harvard Forest, Vermont (looking left and shooting the forest which lies approximately half way between the ground track and Boston. The 1200-hectare site includes a variety of forests and wetlands. Research at the Forest focuses on effects of natural and human disturbances on forest ecosystems).

More... - Chronology...

Home - Browse - Contact
© / Conditions for Use